OpenStudy (anonymous):

solve: (3x+7)/(x+2) + x = 3

myininaya (myininaya):

clear the fraction by multiplying both sides by x+2

OpenStudy (anonymous):

Hint: Find the common denominator

OpenStudy (anonymous):

Do you know how to find common denominators to get a rational fraction?

OpenStudy (anonymous):

no

myininaya (myininaya):

look at what's on that one bottom there which is the x+2 and multiply both sides by x+2 so you don't have to deal with the ugly beast of a fraction

OpenStudy (anonymous):

myininaya as a point - you can do it that way - though I am not sure which way they want you to and if it affects the zeroes that you get out of it.

OpenStudy (anonymous):

i then got 3x+7 + x + 3x+ 6… now what? do i add 3x+x which gives me 3x^2+7=3x+6

myininaya (myininaya):

It looks like you left out the equal sign and forgot to multiply x by also (x+2)

myininaya (myininaya):

\[(x+2) \cdot (\frac{3x+7}{x+2}+x)=(x+2) \cdot 3\] \[(x+2) \cdot \frac{3x+7}{x+2} + (x+2)(x)=3x+6 \]

myininaya (myininaya):

And you were right to cancel the x+2 over x+2 since (x+2)/(x+2)=1

OpenStudy (anonymous):

ok and after that do i just combine like terms and equal it to zero?

myininaya (myininaya):

Well yep you want 0 on one side.

OpenStudy (anonymous):

and once i have zero on one side, i factor out, correct?

myininaya (myininaya):

yep or use quadratic formula

OpenStudy (anonymous):

ok i got x=-1

myininaya (myininaya):

That's right.

OpenStudy (anonymous):

=THANKS!

myininaya (myininaya):

Just make whatever answer you get is in the domain of your expression on both sides which -1 is cool.